• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Armstrong

    Log entry time 10:56:28 on October 5,2009

    Entry number 293256

    keyword=Compton Striptool

    Here is the time history of the Compton. The event with the problem is
    seen at about -30 minutes, where we see the (blue) compton rate and the (magneta) central rate rose, despite the (orange) YPOS not having changed, nor the beam current. You can see our attempts to reset the
    voltage (green and grey lines). You can also see where Mark has set to
    lower voltages. and it stays on).

    FIGURE 1